A solution of cupric sulphate is used at blood donor clinics to test donor blood for low iron. Calculate the concentration of this solution, in mol/L, if 125 g of CuSO4 crystals are dissolved in 1.0 L of water


1
Expert's answer
2021-01-22T06:22:42-0500

c=n/V

n(CuSO4)=m/M=125/160=0.78mol

c=0.78/1=0.78mol/l


Answer: 0.78 mol/l